A metal is tested to have a charge of 697.6C. How many electrons are added to it to acquire the charge?


If two similarly charged metals are brought 1.5m apart in a medium of dielectric constant 2, calculate the force between them.


Is the force attractive or repulsive?

1
Expert's answer
2022-01-12T08:30:00-0500

1.

"N=\frac qe=4.36\cdot 10^{21}."

2.

"F=\frac{kq^2}{\varepsilon r^2}=9.73\cdot 10^{14}~N," repulsive.